# Andrew Gelman

### v1

@@
The ladder of abstraction in statistical graphics [^1] We thank Howard Wainer, Ron Yurko, Dianne Cook, and Alessandra Casella for helpful comments.
Andrew Gelman
(31 May 2023)
1 Overview
@@ 2 Visualizing a mathematical model: basketball shooting
The left plot in Figure 2 shows the path of a ball shot at a $\theta=30^{\circ}$ angle at an initial velocity of $v=15\mbox{m}/\mbox{sec}$. From Newton’s laws of motion and assuming no air resistance, the trajectory is $x(t)=v\cos(\theta)\,t$, $y(t)=a+v\sin(\theta)\,t-0.5gt^{2}$, where $g=9.8\mbox{m}/\mbox{sec}^{2}$ is the gravitational constant. The red line on this graph directly shows the ball’s trajectory. The only way to make it more vivid and less abstract would be an animated presentation that would dynamically show the moving ball.
@@
Plugging in $a=1.70,\theta=30^{\circ},d=10,h=3.05,g=9.8$ into this equation yields $v=12.2$. The resulting trajectory is the blue line in Figure 3. It is helpful to see it in the context of the red curves that show other velocities.
@@
Compared to Figure 3, Figure 4 represents a leap in abstraction: the angle $\theta$ and the initial velocity $v$ are represented by linear coordinates, and the curve is not a trajectory; it si the solution to an equation. Again, in statistics we are so familiar with this sort of graph that we don’t always remember how abstract it is.
@@
The left panel of Figure 5 shows the results. When you are standing close to the basket you need to shoot at a steep angle to reach the hoop; as the distance increases, the optimal angla approaches $45^{\circ}$. The required release velocity increases gradually, at a rate slightly less than linearly with distance. We have lined up the two plots vertically so they can share an $x$-axis and the reader can see how both $\theta_{\rm opt}$ and $v$ vary together.
@@
The right panel of Figure 5 represents the greatest level of abstraction in this series of graphs, and, again, we find it helpful to build up to it rather than simply presenting these results and then expecting readers to follow. That said, we have tried to use graphical cues as much as possible to make plots more intuitive:
@@
The colors climb the ladder of abstraction, from direct mapping of trajectories to the solutions of a series of equations. When presenting the series of graphs we do not draw attention to the colors—that might feel like overkill and distract readers from the content being graphed—but we hope the colors help, in the same way that good design can make a tool easier to use even for users who are unaware of these choices.
@@ 3 Visualizing data: income and voting within and between states
The contrasting patterns for individuals and states motivated us to look at income and voting within states. Figure 7 we shows: Mississippi (the poorest state), Ohio (a middle-income state), and Connecticut (the state with highest median income). You might notice that these are smooth lines rather than the connected points in the left panel of Figure 6. In Figure 7 the lines display summaries of a fitted hierarchical logistic regression model—each of the lines is a part of a logistic curve, with each being on a narrow enough range that the lines are nearly straight.
@@
Indeed, one could take another step of abstraction and summarize each fitted curve by an intercept and slope on the logistic scale (with the income predictor coded as $-2,-1,0,1,2$ so the intercept would be directly interpretable and then make a scatterplot of slope vs. intercept with two-letter state abbreviations. Such a graph could be difficult to explain on its own but could be a natural followup to Figure 7, showing fifty states instead of just three.
@@ 4 Discussion
There is a long history of abstraction in visual representation, going far beyond statistical graphics. Consider, for example, the phrase “the map is not the territory” (Korzybski, 1933) and the satirical vignette of Borges (1946) pointing out the uselessness of a perfectly faithful 1-1 scale map. Figure 11 gives an amusing example discussed by Wainer and Friendly of an early infographic that cleverly conveys information directly through a non-abstract image. But when it comes to displaying mathematical or statistical relationships, we typically take for granted the abstraction of a scatterplot or a graph of $y$ versus $x$.
Modern treatments of data visualization or the grammar of graphics consider graphs not as static drawn objects but as overlays of information, so that a useful plot is built by adding elements to a general structure (Wilkinson, 2005, Wickham, 2016). The ladder of abstraction discussed in the present paper represents a different sort of forward progression: rather than adding information to a graph, we are using increasing levels of generality. Our ideas follow the spirit of the grammer of graphics, however, in considering statistical graphics as a form of communication of quantitative information to the reader, with the specific graph being just an instantiation of this goal. As emphasized by Tufte (1983) and Cleveland (1985), choices in data display can be understood in with respect to the specific goals of communication. In present paper, we emphasize the challenge that useful visualizations can be highly abstract representations of the information being presented, hence our proposal of stepping readers through a series of graphs. We have demonstrated this idea in two different applications of statistical graphics: continuous curves and discrete data.
@@
The present article has three messages. First, some level of abstraction exists in all but the simplest graphs, and we should be aware of this when using graphs to communicate, especially to readers who are not familiar with data graphics. Second, graphical abstraction is not just a difficulty, it also is a useful tool that allows us to understand patterns at a greater level of sophistication, and it can be valuable to see this by adding one level of abstraction at a time rather than trying to jump to a single plot that shows everything. Finally, these principles apply not just to graphs of data but also to graphs of mathematical functions such as fitted models.

### v2

@@
Abstract
@@ 1 Visualizing a mathematical model: basketball shooting
The left plot in Figure 2 shows the path of a ball shot at a $\theta=30^{\circ}$ angle at an initial velocity of $v=15\mbox{m}/\mbox{sec}$. From Newton’s laws of motion and assuming no air resistance, the trajectory is $x(t)=v\cos(\theta)\,t$, $y(t)=a+v\sin(\theta)\,t-0.5gt^{2}$, where $g=9.8\mbox{m}/\mbox{sec}^{2}$ is the gravitational constant. The red line on this graph directly shows the ball’s trajectory. The only way to make it more vivid and less abstract would be an animated presentation that dynamically shows the moving ball.
@@
Plugging $a=1.70,\theta=30^{\circ},d=10,h=3.05,g=9.8$ into this equation yields $v=12.2$. The resulting trajectory is the blue line in Figure 3. It is helpful to see it in the context of the red curves that show other velocities.
@@
Compared to Figure 3, Figure 4 represents a leap in abstraction: the angle $\theta$ and the initial velocity $v$ are represented by linear coordinates, and the curve is not a trajectory; it is the solution to an equation. Again, in statistics we are so familiar with this sort of graph that we don’t always remember how abstract it is.
@@
The left panel of Figure 5 shows the results. When you are standing close to the basket you need to shoot at a steep angle to reach the hoop; as the distance increases, the optimal angle approaches $45^{\circ}$. The required release velocity increases gradually, at a rate slightly less than linearly with distance. We have lined up the two plots vertically so they can share an $x$-axis and the reader can see how both $\theta_{\rm opt}$ and $v$ vary together.
@@
The right panel of Figure 5 represents the greatest level of abstraction in this series of graphs, and, again, we find it helpful to build up to it rather than simply presenting these results and then expecting readers to follow. We also have tried to use graphical cues as much as possible to make plots more intuitive:
@@
The colors climb the ladder of abstraction, from direct mapping of trajectories to the solutions of a series of equations. When presenting the series of graphs we do not draw attention to the colors—that might feel like overkill and distract readers from the content being graphed—but we hope the colors help, in the same way that good design can make a tool more effective even for users who are unaware of these choices.
@@ 2 Visualizing data: income and voting within and between states
The contrasting patterns for individuals and states motivated us to look at income and voting within states. Figure 7 shows Mississippi (the poorest state), Ohio (a middle-income state), and Connecticut (the state with highest median income). You might notice that these are smooth lines rather than the connected points in the left panel of Figure 6. In Figure 7 the lines display summaries of a fitted hierarchical logistic regression model—each of the lines is a part of a logistic curve, with each being on a narrow enough range that the lines are nearly straight.
@@
Indeed, one could take another step of abstraction and summarize each fitted curve by an intercept and slope on the logistic scale (with the income predictor coded as $-2,-1,0,1,2,$ so the intercept would be directly interpretable and then make a scatterplot of slope vs. intercept with two-letter state abbreviations. Such a graph could be difficult to explain on its own but could be a natural followup to Figure 7, showing fifty states instead of just three.
@@ 3 Discussion
There is a long history of abstraction in visual representation, going far beyond statistical graphics. Consider, for example, the phrase “the map is not the territory” (Korzybski, 1933) and the satirical vignette of Borges (1946) pointing out the uselessness of a perfectly faithful 1-1 scale map. Figure 11 gives an amusing example discussed by Wainer and Friendly (2021) of an early infographic that cleverly conveys information directly through a non-abstract image. But when it comes to displaying mathematical or statistical relationships, we typically take for granted the abstractions of scatterplots, time series graphs, bar charts, and so on.
Modern treatments of data visualization or the grammar of graphics consider graphs not as static drawn objects but as overlays of information, so that a useful plot is built by adding elements to a general structure (Wilkinson, 2005, Wickham, 2016). The ladder of abstraction discussed in the present paper represents a different sort of forward progression: rather than adding information to a graph, we are using increasing levels of generality. Our ideas follow the spirit of the grammer of graphics, however, in considering statistical graphics as a form of communication of quantitative information to the reader, with the specific graph being just an instantiation of this goal. As emphasized by Tufte (1983) and Cleveland (1985), choices in data display can be understood with respect to the specific goals of communication. In present paper, we emphasize the challenge that useful visualizations can be highly abstract representations of the information being presented, hence our proposal of stepping readers through a series of graphs. Here we have demonstrated this idea in two different applications of statistical graphics: continuous curves and discrete data.
@@
The present article has three messages. First, some level of abstraction exists in all but the simplest graphs, and we should be aware of this when using graphs to communicate, especially to readers who are not familiar with data graphics. Second, graphical abstraction is not just a difficulty; it also is a useful tool that allows us to understand patterns at a greater level of sophistication, and it can be valuable to see this by adding one level of abstraction at a time rather than trying to jump to a single plot that shows everything. Finally, these principles apply not just to graphs of data but also to graphs of mathematical functions such as fitted models.